# When and Where do Data Poisons Attack Textual Inversion?

### v2

@@ 5.2.3 Masking Ablation
In Appendix J, we experiment with dilating the mask size to incorporate additional context in loss backpropagation. We find that relatively small amounts of mask dilation (i.e., $16$ pixels for $512$x512512512512 masks) can improve LM for TI.
@@ 5.2.4 Safe-Zone Training Performance
Figure 8: Generated images after TI on NovelConcepts10 BlueElephant poisoned by ADM+. Generation prompt: “a $R*$ on the beach"
Our implementation of SZT is highly customizable, empowering the user to control JPEG compression quality, timestep restriction method/shape, and loss mask size/shape. In our ablations, we find that combining JPEG compression (quality$=25$), timestep restriction $t\geq 500$, and LM with $16$ pixels of mask dilation performs best. We defer further discussion on SZT configurations to Appendix J and encourage users to explore additional settings.
@@ Appendix B Noise-Prediction Distribution
$p(\epsilon,z_{t}|t)=\mathcal{N}\left(\begin{bmatrix}0\\ \sqrt{\overline{\alpha}_{t}}\mu_{z}\end{bmatrix},\\ \begin{bmatrix}I&\sqrt{1-\overline{\alpha}_{t}}I\\ \sqrt{1-\overline{\alpha}_{t}}I&\Sigma_{zt}\end{bmatrix}\right),$ (9)
@@
Figure 12: Histograms of pixel-space perturbations (relative to clean images) for poisoned NovelConcept10 images without (red) and with (green) JPEG compression. Input images are shifted and scaled to $[-1,1]$
Fig. 10 is an expanded version of Fig. 4 that includes plots for the magnitudes of the predicted noise. As discussed in Sec. 4.2, the model predicts noise near 00 at $t=0$ and perfectly predicts noise at $t=T$. The magnitude of the predicted noise grows from near 00 at $t=0$ to a maximum of $128$ at $t=T$ (which aligns with $E[||\epsilon||_{2}]=\sqrt{tr(I)}=128$ when $\epsilon$ is a 4x64x64 vector as in Stable Diffusion). The median loss monotonically decreases from a maximum at $t=0$ to 00 at $t=T$. Additionally, the width of the loss distribution at each timestep indicates that the variance of the noise-prediction error is indeed a maximum at $t=0$ and diminishes to 00 at $t=T$.
@@ Appendix C Textual Inversion Gradients
In Fig. 11 we display the $l^{2}$ magnitude of the loss gradient on the trainable text embedding as a function of timestep $t$ throughout TI training. For each concept in the NovelConcepts10 dataset, we extract checkpoints every 1000 steps during TI training and calculate the expected gradient magnitude for each timestep then average the magnitudes across concepts. Aligning with our findings from SSMs (Sec. 4.1) and timestep learning bias (Sec. 4.2), all gradient magnitudes are biased towards lower-middle timesteps, with near-00 gradient magnitudes at high timesteps for all datasets. The ADM- poison, which seeks to minimize DM loss, mitigates the low-timestep bias. The EA poison, which optimizes poisons on the LDM encoder and avoids the DM entirely during optimization, still permits a bias towards lower timesteps during TI training. This is likely because EA focuses on perturbing $z_{0}$, thereby adversarially affecting high signal-to-noise $z_{t}$ latents (i.e., those at lower-middle timesteps).
@@ Appendix D JPEG Compression Analysis
In Fig. 12, we extend the analysis performed for Fig. 7 in Sec. 4.4 to all poisons. Across all poisons, we find that JPEG compression has the same effect: the poison perturbation distribution is converted from a bimodal distribution with modes that skew towards the perturbation limits (±$\pm\kappa$) to a bell-curve distribution centered at or near 00.
To verify claim (2) from Sec 4.4, we also investigate the impact of JPEG compression on the LDM latent space. Fig. 13 displays the radially averaged power spectrum density curves for the mean latent encodings of Clean, ADM+, ADM-, and EA images with and without JPEG compression, averaged across all 50 images in the NovelConcepts10 dataset. Before JPEG compression, the power curve for ADM+ latents is consistently higher than that of clean latents at high frequencies, whereas the powers for ADM- and EA latents are consistently lower. However, after JPEG compression (in pixel space), the power spectra of all images are centralized, lying closer to the power spectra of clean images. This result suggests that JPEG compression tends to “standardize" poisoned images, forcing their latents to conform to the power spectra expected by LDMs.
@@ Appendix H Timestep Range Ablation
Figure 15: Probability curves for timestep sampling, displaying $tanh$ distributions (top) and power distributions (bottom). Distribution parameters for each curve are given in the legends.
@@ Appendix I Masking Ablation
Figure 16: Average $R_{n,M_{z}}$ for clean and poisoned versions of NovelConcepts10 throughout TI training. “wLM" denotes TI with loss masking.
@@
In initial experiments with SZT, we found that combining T600 and LM tended to perform below expectations. Although the performance of T600+LM was typically better than established defenses like Regen and AdvClean, it underperformed other SZT ablations like JPEG+T600 or JPEG+LM. We hypothesize that strong restrictions in both time (i.e., T600) and space (i.e., LM) may be too restrictive and may hinder concept learning during TI training. Therefore, we evaluated additional configurations of SZT that ease temporal and spatial restrictions when used in combination. In particular, we investigate lower timestep threshold (e.g., $t\geq 500$, denoted as “T500") combined with dilated concept masks. To implement mask dilations for LM, we apply the ImageFilter.MaxFilter method from PIL with $8$, $16$, or $24$ pixels of dilation to the $512$x512512512512 binary masks, then rescale to $64$x64646464 for latent space. Intuitively, applying mask dilation includes extra background information outside of the novel concept region during loss backpropagation.
Tables 12, 13, and 14 display the DINOv2 similarity, FID, and CLIP Score for various combinations of T500, T600, and dilated masks (denoted “LM-D08", “LM-D16", and “LM-D24"). Of the various LM configurations, $16$ pixels of dilation (LM-D16) demonstrates the highest robustness to poisons. Furthermore, when combining T500 or T600 with LM, using T500+LM-D16 gives the best performance, supporting our “too restrictive" hypothesis above. Our final implementation of SZT uses JPEG preprocessing with T500+LM-D16 and further improves poison defense beyond all other ablations.
@@
For emphasis, we plot the DINOv2 Similarity versus CLIP Score values for “Psn Avg" on CustomConcept101 for all defenses in Fig. 17. As discussed in Sec. 5.2.4, Regen, PDMPure, and AdvClean all offer minor improvements in poison performance. We note the that despite a drastic improvement in DINOv2 Similarity, JPEG is limited in its prompt fidelity (measured by CLIP Score). This aligns with qualitative observations of its limited concept learning from Fig. 8. SZT improves DINOv2 Similarity and CLIP Score beyond all existing defenses, and most ablations of SZT also perform well. We note that all ablations of SZT can beat existing defenses in at least one metric.
Figure 17: DINOv2 Similarity versus CLIP Score for all defenses on CustomConcept101. Values are from the “Psn Avg" rows in Tables 1 and 16.
@@
Figure 18: Images generated after TI on various concepts from CustomConcept101 poisoned by ADM+. Prompt: “a $R*$ in the snow"
Figure 19: Images generated after TI on things-bottle1 (CustomConcept101) for various poisons. Prompt: “a $R*$ in the snow"
Figure 20: Images generated after TI on various concepts from NovelConcepts10 poisoned by ADM+. Prompt: “a $R*$ on the beach"
Figure 21: Images generated after TI on FishDoll (NovelConcepts10) for various poisons. Prompt: “a $R*$ on the beach"
@@
This experiment approaches a “black-box" poisoning scenario, as we craft all poisons using SD15 and then test defenses on different Stable Diffusion versions. This is not a major concern, as multiple prior works have demonstrated the strong transferability of poisons across Stable Diffusion models [78, 37]. In multiple tables for SD15rf, SD21, and SD21v results below, the “Nominal" defense column indeed shows that all poisons are effective. Network-based defenses (Regen and PDMPure) shown here still use the same models for purification.
@@
Poison and defense performance on SD21 in Tables 26 and 27 is generally similar to that of SD15, which is sensible given their similar architectures, training data, and training objectives. The trends for SD21v in Tables 28 and 29 are generally similar to those of SD15 and SD21, though the effect of poisoning (i.e., the difference between “Clean" and “Psn Avg" values) is diminished relative to SD15 and SD21. Even so, SZT and its ablations are still effective for improving generation quality on poisoned data. In Tables 30 and 31, SDXL demonstrates the largest rift in performance trends relative to SD15, being unaffected by most poisons. Though SZT and its ablations can improve performance on poison data they are generally not necessary. We are uncertain of the exact cause of the low degree of poison transferability from SD15 to SDXL.
@@
In our implementation of LoRA, we use rank $4$ adapters with a target prompt that includes “*" as a dummy token to signify LoRA usage and trained for $1000$ steps. We did not find any improvement in generative quality when using the prior preservation loss or when training text encoder weights, and thus our implementation does not include these methods. We report LoRA results in Tables 32, 33, and 34. We find that poisons are moderately effective against LoRA, with EA being the most severe. Existing defenses like PDMPure and AdvClean slightly improve defense against poisons but Regen generally reduces generative quality. SZT and its ablations outperform existing defenses to achieve clean-level generative quality for poisoned images. We note that as LoRA does not provide any finetuned text tokens, the CLIP Score becomes a metric for background fidelity; the CLIP text encoder will ignore the dummy token (“*") and instead focus on context words.

### v4

@@ 5.2.3 Masking Ablation
In Appendix J, we experiment with dilating the mask size to incorporate additional context in loss backpropagation. We find that relatively small amounts of mask dilation (i.e., $16$ pixels for 512512x512512 masks) can improve LM for TI.
@@ 5.2.4 Safe-Zone Training Performance
Figure 8: Generated images after TI on NovelConcepts10 BlueElephant poisoned by ADM+. Generation prompt: “a RR* on the beach”
Our implementation of SZT is highly customizable, empowering the user to control JPEG compression quality, timestep restriction method/shape, and loss mask size/shape. In our ablations, we find that combining JPEG compression (quality=25=25), timestep restriction $t\geq 500$, and LM with $16$ pixels of mask dilation performs best. We defer further discussion on SZT configurations to Appendix J and encourage users to explore additional settings.
@@ Appendix B Noise-Prediction Distribution
[…]
@@
Figure 12: Histograms of pixel-space perturbations (relative to clean images) for poisoned NovelConcept10 images without (red) and with (green) JPEG compression. Input images are shifted and scaled to [1,1][-1,1]
[…]
@@ Appendix C Textual Inversion Gradients
In Fig. 11 we display the $l^{2}$ magnitude of the loss gradient on the trainable text embedding as a function of timestep $t$ throughout TI training. For each concept in the NovelConcepts10 dataset, we extract checkpoints every 1000 steps during TI training and calculate the expected gradient magnitude for each timestep then average the magnitudes across concepts. Aligning with our findings from SSMs (Sec. 4.1) and timestep learning bias (Sec. 4.2), all gradient magnitudes are biased towards lower-middle timesteps, with near-0 gradient magnitudes at high timesteps for all datasets. The ADM- poison, which seeks to minimize DM loss, mitigates the low-timestep bias. The EA poison, which optimizes poisons on the LDM encoder and avoids the DM entirely during optimization, still permits a bias towards lower timesteps during TI training. This is likely because EA focuses on perturbing $z_{0}$, thereby adversarially affecting high signal-to-noise $z_{t}$ latents (i.e., those at lower-middle timesteps).
@@ Appendix D JPEG Compression Analysis
In Fig. 12, we extend the analysis performed for Fig. 7 in Sec. 4.4 to all poisons. Across all poisons, we find that JPEG compression has the same effect: the poison perturbation distribution is converted from a bimodal distribution with modes that skew towards the perturbation limits ($\pm\kappa$) to a bell-curve distribution centered at or near 0.
To verify claim (2) from Sec 4.4, we also investigate the impact of JPEG compression on the LDM latent space. Fig. 13 displays the radially averaged power spectrum density curves for the mean latent encodings of Clean, ADM+, ADM-, and EA images with and without JPEG compression, averaged across all 50 images in the NovelConcepts10 dataset. Before JPEG compression, the power curve for ADM+ latents is consistently higher than that of clean latents at high frequencies, whereas the powers for ADM- and EA latents are consistently lower. However, after JPEG compression (in pixel space), the power spectra of all images are centralized, lying closer to the power spectra of clean images. This result suggests that JPEG compression tends to “standardize” poisoned images, forcing their latents to conform to the power spectra expected by LDMs.
@@ Appendix H Timestep Range Ablation
Figure 15: Probability curves for timestep sampling, displaying tanhtanh distributions (top) and power distributions (bottom). Distribution parameters for each curve are given in the legends.
@@ Appendix I Masking Ablation
Figure 16: Average $R_{n,M_{z}}$ for clean and poisoned versions of NovelConcepts10 throughout TI training. “wLM” denotes TI with loss masking.
@@
In initial experiments with SZT, we found that combining T600 and LM tended to perform below expectations. Although the performance of T600+LM was typically better than established defenses like Regen and AdvClean, it underperformed other SZT ablations like JPEG+T600 or JPEG+LM. We hypothesize that strong restrictions in both time (i.e., T600) and space (i.e., LM) may be too restrictive and may hinder concept learning during TI training. Therefore, we evaluated additional configurations of SZT that ease temporal and spatial restrictions when used in combination. In particular, we investigate lower timestep threshold (e.g., $t\geq 500$, denoted as “T500”) combined with dilated concept masks. To implement mask dilations for LM, we apply the ImageFilter.MaxFilter method from PIL with $8$, $16$, or $24$ pixels of dilation to the 512512x512512 binary masks, then rescale to 6464x6464 for latent space. Intuitively, applying mask dilation includes extra background information outside of the novel concept region during loss backpropagation.
Tables 12, 13, and 14 display the DINOv2 similarity, FID, and CLIP Score for various combinations of T500, T600, and dilated masks (denoted “LM-D08”, “LM-D16”, and “LM-D24”). Of the various LM configurations, $16$ pixels of dilation (LM-D16) demonstrates the highest robustness to poisons. Furthermore, when combining T500 or T600 with LM, using T500+LM-D16 gives the best performance, supporting our “too restrictive” hypothesis above. Our final implementation of SZT uses JPEG preprocessing with T500+LM-D16 and further improves poison defense beyond all other ablations.
@@
For emphasis, we plot the DINOv2 Similarity versus CLIP Score values for “Psn Avg” on CustomConcept101 for all defenses in Fig. 17. As discussed in Sec. 5.2.4, Regen, PDMPure, and AdvClean all offer minor improvements in poison performance. We note the that despite a drastic improvement in DINOv2 Similarity, JPEG is limited in its prompt fidelity (measured by CLIP Score). This aligns with qualitative observations of its limited concept learning from Fig. 8. SZT improves DINOv2 Similarity and CLIP Score beyond all existing defenses, and most ablations of SZT also perform well. We note that all ablations of SZT can beat existing defenses in at least one metric.
Figure 17: DINOv2 Similarity versus CLIP Score for all defenses on CustomConcept101. Values are from the “Psn Avg” rows in Tables 1 and 16.
@@
Figure 18: Images generated after TI on various concepts from CustomConcept101 poisoned by ADM+. Prompt: “a RR* in the snow”
Figure 19: Images generated after TI on things-bottle1 (CustomConcept101) for various poisons. Prompt: “a RR* in the snow”
Figure 20: Images generated after TI on various concepts from NovelConcepts10 poisoned by ADM+. Prompt: “a RR* on the beach”
Figure 21: Images generated after TI on FishDoll (NovelConcepts10) for various poisons. Prompt: “a RR* on the beach”
@@
This experiment approaches a “black-box” poisoning scenario, as we craft all poisons using SD15 and then test defenses on different Stable Diffusion versions. This is not a major concern, as multiple prior works have demonstrated the strong transferability of poisons across Stable Diffusion models [78, 37]. In multiple tables for SD15rf, SD21, and SD21v results below, the “Nominal” defense column indeed shows that all poisons are effective. Network-based defenses (Regen and PDMPure) shown here still use the same models for purification.
@@
Poison and defense performance on SD21 in Tables 26 and 27 is generally similar to that of SD15, which is sensible given their similar architectures, training data, and training objectives. The trends for SD21v in Tables 28 and 29 are generally similar to those of SD15 and SD21, though the effect of poisoning (i.e., the difference between “Clean” and “Psn Avg” values) is diminished relative to SD15 and SD21. Even so, SZT and its ablations are still effective for improving generation quality on poisoned data. In Tables 30 and 31, SDXL demonstrates the largest rift in performance trends relative to SD15, being unaffected by most poisons. Though SZT and its ablations can improve performance on poison data they are generally not necessary. We are uncertain of the exact cause of the low degree of poison transferability from SD15 to SDXL.
@@
In our implementation of LoRA, we use rank $4$ adapters with a target prompt that includes “*” as a dummy token to signify LoRA usage and trained for $1000$ steps. We did not find any improvement in generative quality when using the prior preservation loss or when training text encoder weights, and thus our implementation does not include these methods. We report LoRA results in Tables 32, 33, and 34. We find that poisons are moderately effective against LoRA, with EA being the most severe. Existing defenses like PDMPure and AdvClean slightly improve defense against poisons but Regen generally reduces generative quality. SZT and its ablations outperform existing defenses to achieve clean-level generative quality for poisoned images. We note that as LoRA does not provide any finetuned text tokens, the CLIP Score becomes a metric for background fidelity; the CLIP text encoder will ignore the dummy token (“*”) and instead focus on context words.